# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions

### v1

@@
StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions
Nicholas Kraabel
Department of Civil and Environmental Engineering
Pennsylvania State University
&Jiangtao Liu11footnotemark: 1
Department of Civil and Environmental Engineering
Pennsylvania State University
&Yuchen Bian
Amazon Web Services
&Daniel Kifer
Department of Computer Science & Engineering
Pennsylvania State University
&Chaopeng Shen
Department of Civil and Environmental Engineering
Pennsylvania State University
Equal contribution.Corresponding author. cshen@engr.psu.edu
Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy. Traditional impact models, whether process-based, statistical, or machine learning, struggle with spatial generalization due to limited observations and concept drift. Recently proposed vision foundation models trained on satellite imagery demand massive compute and are ill-suited for dynamic land-surface prediction. We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions. StefaLand improves predictions on three tasks and four datasets: streamflow, soil moisture, and soil composition, compared to prior state-of-the-art. Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications. The model builds on a masked autoencoder backbone that learns deep joint representations of landscape attributes, with a location-aware architecture fusing static and time-series inputs, attribute-based representations that drastically reduce compute, and residual fine-tuning adapters that enhance transfer. While inspired by prior methods, their alignment with geoscience and integration in one model enables robust performance on dynamic land–surface tasks. StefaLand can be pretrained and finetuned on academic compute yet outperforms state-of-the-art baselines and even fine-tuned vision foundation models. To our knowledge, this is the first geoscience land-surface foundation model that demonstrably improves dynamic land-surface interaction predictions and supports diverse downstream applications.
@@ 1 Introduction
Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards Ebi et al. (2021), IPCC (2021). To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content. Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions Dorigo et al. (2013). Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively. Soil composition (sand, silt, clay fractions) governs infiltration capacity and root-zone storage, while slope–soil-vegetation interactions directly influence landslide hazards. Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change. Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
@@
Traditionally, these tasks were undertaken by physically-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest Li et al. (2015). In recent years, there has been a proliferation of data-driven machine learning (ML) models Solomatine and Ostfeld (2008). These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs. However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks. Foundation models, which are trained on broad tasks to grasp the joint data distribution, have just started to emerge in geosciences, e.g., Terramind, Prithvi and Aurora. However, they were built mainly on satellite images for landcover-identification tasks Jakubik et al. (2023; 2025), Schmude et al. (2024). There has been a notable absence of foundation models built for dynamical land surface modeling, and a lack of effort to leverage valuable temporal datasets and ground-based observations Xie et al. (2023).
@@
A Grand challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data. Due to the cost of installing instruments and varying policies on data sharing, data are often only available in high density in certain regions of some developed nations. For example, streamflow gauge data are available at high density in United States, Europe, Australia and Japan, but are scanty in Africa, South America, and the rest of the Asia Global Runoff Data Centre (2020). Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements. Satellite data have too coarse resolution and too large uncertainty levels compared to in-situ measurements. As quantified in many studies Feng et al. (2023), a deep network trained using data from some regions can face substantial performance degradation when applied in data-scarce regions. This is partly because there are not enough sites in space to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies among different regions (concept drift). This means ML land surface predictions may be unusable in more than 70% of the land. This limitation is particularly acute in data-scarce regions that are often most vulnerable to climate impacts. Any method that can reliably improve spatial generalization can be a significant and rare advance to greatly democratizes information for stakeholders worldwide.
@@
Related Work: In hydrologic and ecosystem predictions, long short-term memory (LSTM) networks Hochreiter and Schmidhuber (1997) remain the dominant architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well. Attempts to adopt transformers, so successful in natural language processing, have generally underperformed in hydrologic regression tasks Xue et al. (2023), Liu et al. (2024), with evidence of overfitting on continuous signals Zeng et al. (2022). Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results and even offer advantages in autoregressive forecasting and extreme event prediction Wen et al. (2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
@@
Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class Hrachowitz et al. (2013), Yang et al. (2023). Such an expert-derived design represents a crude practice of unsupervised learning that indicate the importance of understanding the joint data distribution. However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches. Foundation models offer a promising approach to address these spatial generalization challenges. By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains Zhang et al. (2024).
@@
Existing geoscience foundation models such as TerraMind, Prithvi, Aurora, and AlphaEarth have largely focused on satellite imagery Hsu et al. (2024), which, while providing global coverage, may not capture the temporal dynamics and physical processes most relevant to land surface predictions. For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation Jakubik et al. (2023). It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" Jakubik et al. (2025). However, many critical variables regarding soil and the subsurface are not directly observable from space. In addition, satellite images are noisy and a large portion of the data is redundant and irrelevant across repeated revisits. Therefore, it is unclear whether they are of high value to land-surface dynamical prediction tasks.
@@
Our contributions: We present Spatiotemporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction. StefaLand improves predictions on streamflow, soil moisture, and soil composition compared to state-of-the-art baselines. StefaLand further demonstrates strong spatial generalization across diverse landscapes and data-scarce regions. StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) is intentionally aligned with physical processes, drastically reducing compute requirements while retaining global coverage, making it accessible to academic researchers with modest resources. Pretraining requires only about 720 GPU hours, which makes it feasible within academic budgets. The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge. Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models, purpose-built for hydrologic and land–surface applications.
@@ 2 Methods
Dynamic land–surface prediction requires integrating heterogeneous information: landscape attributes such as topography, soils, and geology, together with dynamic meteorological forcings like precipitation and temperature. Capturing the joint influence of these variables is essential for spatial generalization but challenging with standard architectures. StefaLand addresses this gap through an attribute-based transformer framework that unifies static and time-varying signals in a shared representation space.
2.1 Model Architecture Overview
StefaLand processes both dynamic time series and static attributes through a unified encoder-decoder architecture, and is trained as a attribute-group-based masked autoencoder (Figure D1). The model embeds multivariate inputs, applies strategic masking to force the learning of cross-domain dependencies, and reconstructs the masked portions to generate deep representations. Our location-aware transformer encoder captures spatial relationships while positional encodings maintain temporal context. The complete mathematical formulation of our model architecture is provided in Appendix A. As well as a visual representation of the model is provide in appendix D.
2.2 Pretraining Setup
Our pretraining approach uses Cross-Variable Group Masking (CVGM), which forces the model to capture interactions among groups of landscape variables (e.g., terrain, climate, soil, vegetation) rather than treating them independently. This encourages the learning of physical relationships between different environmental drivers. The objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector. During pretraining, the dataset is stored in shards rather than fully loaded into memory. At the start of each epoch, we randomly reindex these shards, which effectively reshuffles the available samples. This lazy-loading strategy avoids materializing the entire corpus in memory while still ensuring high sample diversity across epochs.
The pretraining dataset is a derived global attribute dataset spanning 4,229 locations over 40 years. A complete list of variables and their sources is provided in Appendix C. The accompanying code release will enable reconstruction of this dataset. Hyperparameter settings and configurations for pretraining are also listed in the appendix above.
2.3 Transformer-based Direct Prediction Models
We developed several finetuned transformer models that combine StefaLand’s base structure with new units to directly predict streamflow and soil moisture. StefaLand with residual connection (resConn) is our main proposed architecture, integrating pretrained StefaLand transformers with an LSTM decoder through a residual pathway (Figure D1). This architecture enables iterative integration of pretrained transformer features throughout the decoding process, combining frozen transformer embeddings with forcings via a convolutional neural network before merging with the decoder through skip connections. This enables effective integration of pretrained features with task-specific requirements. Detailed adapter architectures are described in Appendix A.
For comparisons, "StefaLand without resConn" utilizes transformer outputs only once through a simple adapter (Figure D2), and a comparison with this model highlights the value of the resConn structure. "StefaLand Transformer" represents the base StefaLand model trained from scratch with the finetuning dataset. "StefaLand Ablation - resConn" utilizes the same structure as "StefaLand - resConn", but randomize the weights of StefaLand instead of using the pretrained weights. Comparing with these two models demonstrates the value of pretraining. As discussed below, we also implemented Terramind with resConn, which uses IBM’s Terramind foundational model together with our residual connection architecture, to contrast the effectiveness of Terramind vs StefaLand pretraining. Of course, supervised LSTM serves as a widely employed and highly-competitive baseline.
2.4 Physics-Based Integration
To leverage domain knowledge, we also implemented physics-based configurations that combine our data-driven approach with process understanding. For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model, a conceptual hydrologic model with state variables like snow storage, soil water, and subsurface storage. In this setup, the neural network only provides parameters for the process-based backbone, which then outputs a range of interpretable variables, and they are trained together in an end-to-end fashion.
Our physics-based approaches include LSTM-HBV1.1, a baseline configuration where LSTM outputs parameters for the HBV model; StefaLand-resConn HBV1.1; and StefaLand-no resConn HBV1.1, which integrates transformer features without residual connections. These physics-based approaches maintain interpretability while leveraging improved representations from StefaLand. Detailed implementations are provided in Appendix B.

### v2

# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions

@@
Abstract
Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy. Traditional impact models, whether process-based, statistical, or machine learning, struggle with spatial generalization due to limited observations and concept drift. Recently proposed vision foundation models trained on satellite imagery demand massive compute and are ill-suited for dynamic land-surface prediction. We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions. StefaLand improves predictions on four tasks and five datasets: streamflow, soil moisture, soil composition, and landslide detection compared to prior state-of-the-art. Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications. The model builds on a masked autoencoder backbone that learns deep joint representations of landscape attributes, with a location-aware architecture fusing static and time-series inputs, attribute-based representations that drastically reduce compute, and residual fine-tuning adapters that enhance transfer. While inspired by prior methods, their alignment with geoscience and integration in one model enables robust performance on dynamic land–surface tasks. StefaLand can be pretrained and finetuned on academic compute yet outperforms state-of-the-art baselines and even fine-tuned vision foundation models. To our knowledge, this is the first geoscience land-surface foundation model that demonstrably improves dynamic land-surface interaction predictions and supports diverse downstream applications.
@@ 1 Introduction
Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021, IPCC, 2021). To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content. Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions (Dorigo et al., 2013a). Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively. Soil composition (sand, silt, clay fractions) governs infiltration capacity and root-zone storage, while slope–soil-vegetation interactions directly influence landslide hazards. Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change. Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
@@
Traditionally, these tasks were undertaken by physics-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest (Li et al., 2015). In recent years, there has been a proliferation of data-driven machine learning (ML) models (Solomatine and Ostfeld, 2008). These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs. However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks. A particularly notable gap is that currently established geoscience foundation models have largely been trained on satellite imagery for landcover-identification tasks (Jakubik et al., 2023; 2025, Schmude et al., 2024), which limits their ability to capture dynamic land–atmosphere interactions. As a result, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
@@
A grand challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data. Satellite data are often coarse in resolution and uncertain compared to in-situ measurements. SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010). However, in-situ data, due to the cost of installing instruments and varying policies on data sharing, is only available in high density in certain regions of some developed nations. For example, streamflow gauge data are abundant in the United States, Europe, Australia and Japan, but remain sparse in Africa, South America, and much of Asia (Global Runoff Data Centre, 2020). Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements. As quantified in many studies (Feng et al., 2023), a deep network trained on data from some regions can face substantial performance degradation when applied in data-scarce regions. This occurs partly because there are not enough sites to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies across regions (concept drift). While such limitations hinder traditional supervised ML models, foundation models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
@@
Related Work: In hydrologic and ecosystem predictions, supervised long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) remain a highly popular architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well(Kratzert et al., 2018). Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023, Liu et al., 2024), with evidence of overfitting on continuous signals (Zeng et al., 2022). Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
@@
Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013, Yang et al., 2023). Such an expert-derived design represents a crude practice of unsupervised learning that indicates the importance of understanding the joint data distribution. However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches. Foundation models offer a promising approach to address these spatial generalization challenges. By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains (Zhang et al., 2024).
@@
Existing geoscience foundation models, e.g., TerraMind (Jakubik et al., 2025) and Prithvi (Hsu et al., 2024) from IBM, and Aurora (Bodnar et al., 2025) from Microsoft, have largely focused on satellite imagery, which may not capture the temporal dynamics and physical processes most relevant to land surface predictions. For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation (Jakubik et al., 2023). It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" (Jakubik et al., 2025). However, many critical variables regarding soil and the subsurface are not directly observable from space. In addition, satellite images are noisy and a large portion of the data is redundant across repeated revisits or irrelevant for hydrologic processes. Therefore, it is unclear whether they are of high relevance to land-surface dynamical prediction tasks.
@@
Our contributions: We present Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction. StefaLand improves predictions on streamflow, soil moisture, soil composition, and landslide susceptibility compared to state-of-the-art supervised learning baselines as well as finetuned satellite-image-trained earth foundation models. Especially, it shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks. StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) incorporates a variety of ground-based measurement data, emphasizes relevant land-surface physical processes, drastically reduces compute requirements while retaining global coverage, making it accessible to researchers with modest resources. Pretraining our model required only about 720 V100 GPU hours (could be shorter with more advanced GPUs). The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge. Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models.
@@ 2 Methods
Dynamic land–surface prediction requires combining heterogeneous information: static landscape attributes such as topography, soils, vegetation, and geology, together with dynamic forcings such as precipitation and temperature. StefaLand addresses this challenge with a transformer-based masked autoencoder that jointly embeds static and dynamic variables, pretrained with a cross-variable masking strategy, and then adapted for specific prediction tasks with task-specific heads.
2.1 StefaLand Transformer Architecture
Embedding.
Each dynamic input variable $c$ at time $t$ is independently embedded with a nonlinear two-layer projection:
$z_{t,c}=\text{GELU}(x_{t,c}W_{1,c}+b_{1,c})W_{2,c}+b_{2,c},$ (1)
where $W_{1,c}\in\mathbb{R}^{1\times 64}$ and $W_{2,c}\in\mathbb{R}^{64\times 256}$. Summing across all $C$ dynamic variables yields a per-step embedding:
$z_{t}=\sum_{c=1}^{C}z_{t,c}.$ (2)
Static attributes $s_{i}$ are embedded similarly:
$z_{\text{static},i}=\text{GELU}(s_{i}W_{1,i}+b_{1,i})W_{2,i}+b_{2,i},$ (3)
and concatenated as a special static token, producing the joint sequence:
$Z=[z_{1};z_{2};\ldots;z_{T};z_{\text{static}}].$ (4)
Cross-Variable Group Masking (CVGM).
Let $G=\{g_{1},\ldots,g_{K}\}$ be groups of related variables. During pretraining, a temporal window $\tau,\tau+\ell)$ is sampled, and entire groups are stochastically masked:
$m_{k}\sim\text{Bernoulli}(p_{\text{mask}}).$ (5)
For all $c\in g_{k}$ with $m_{k}=1$, the embedding is replaced by a learned mask vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$. This forces reconstructions to rely on cross-variable dependencies rather than treating drivers independently.
Transformer encoder.
After adding learnable positional encodings $P$, we obtain:
$\tilde{Z}=Z+P.$ (6)
This sequence is processed by $N$ Transformer blocks:
$A^{(\ell)}$ $=\text{MHA}(H^{(\ell-1)}),$ (7)
$\tilde{H}^{(\ell)}$ $=\text{LayerNorm}(H^{(\ell-1)}+A^{(\ell)}),$ (8)
[…]
$H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)}).$ (10)
Decoder and loss.
The final hidden states are passed through a bidirectional LSTM to refine local temporal continuity:
$U=\text{LSTM}(H^{(N)}W_{\text{proj}}+b_{\text{proj}}).$ (11)
Outputs corresponding to masked groups are projected back to the original variable dimensions. The reconstruction loss is
$\mathcal{L}=\sum_{c\in\mathcal{M}}w_{c}\frac{\|\hat{x}_{c}-x_{c}\|_{2}^{2}}{\sigma_{c}^{2}},$ (12)
where $\mathcal{M}$ is the set of masked variables, $\sigma_{c}$ is the standard deviation of variable $c$, and $w_{c}$ is a learnable weight.
2.2 Pretraining Details
The pretraining dataset is a derived global attribute dataset spanning $\sim$8,000 locations (basins) over 40 years. Variables were chosen to represent the key controls on fluxes of water, energy, momentum, sediment, and nutrients. A complete list of variables, their group assignments, and their sources is provided in Appendix C.
The CVGM is designed so variables with reciprocal or bidirectional causality are masked together, preventing them from acting as predictors for each other. Most groupings are straightforward, such as masking silt and clay percentages together. One exception is soil depth, which we classify as a terrain attribute since it is strongly tied to topographic derivations. By masking and reconstructing at the group level, the model is encouraged to capture cross-domain interactions, such as the coupling between soil texture and climate seasonality. While the grouped-masking strategy has sometimes been employed in multimodality AI models, we did not find it applied in earth foundation models like Terramind, Prithvi, or Aurora, which instead use random dropout, spatial patches, or future periods for masking. Our objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector.
2.3 Finetuning for Prediction Tasks
Our primary finetuning model, StefaLand-resConn, integrates pretrained embeddings with raw forcings through a residual pathway (Figure D1). Let $E_{t}$ denote the transformer embedding at time $t$, and $x_{t}$ the raw forcings. We compute:
$r_{t}$ $=f_{\text{conv+linear}}(x_{t}),$ (13)
$h_{t}$ $=\text{LSTM}(E_{t}+r_{t}),$ (14)
$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o}.$ (15)
Skip connections propagate $E_{t}$ and $r_{t}$ into intermediate layers, allowing general knowledge (from $E_{t}$) to be iteratively refined with task-specific signals ($r_{t}$). This design strengthens spatial generalization while adapting to local variability. To see a visual representation of this D1.